A company had 8.28 million shares in issue at the start of the year and made no new issue of
shares during the year ended 31st December 2004, but on that date there were outstanding
options to purchase 920,000 ordinary $1 shares at $1.7 per share. The average fair value of
ordinary shares was $1.80. Earnings for the year ended 31st December 2004 were $2,208,000.
Required
Calculate the fully DEPS for the year ended 31st December, 2004
1
Expert's answer
2020-12-28T08:22:01-0500

Given is the following information

Earnings for the year ended 31st December 2004 = $2,208,000

Number of share basic = 8.28 million

Number of option share =

"=920000\times(\frac{1.80-1.70}{1.80})"

=51,111

Total number of weighted outstanding share = 8,331,111

Diluted earning per share = 2,208,000/8,331,111

=0.26